OpenStudy (anonymous):

f(x)=x^3 + 2x^2 + x solve for x

OpenStudy (callisto):

x^3 + 2x^2 + x = 0 Factorize the left side x(x^2+2x+1) =0 Use the idenity (a+b)^2 = a^2 + 2ab + b^2 to factorize x^2+2x+1 finally put each factor =0 and solve x